OpenStudy (anonymous):

Use addition or substitution to find the value of y for this set of equations. x + y = 14 x = 6y

OpenStudy (zpupster):

Substitution first: to the 1st equation lets subtract x from both sides x-x+y = 14 - x y=14 - x now we know what y is = to lets substitute that in the 2nd equation x = 6(14 - x) multiply 6 by all the terms in the bracket x = 84 - 6x add 6x to both sides x + 6x = 84 - 6x + 6x combine our terms 7x = 84 divide 7 to both sides 7x/7 = 84 /7 x = 12 we know from before y = 14 - x plug in x y = 14 - 12 y= 2

OpenStudy (anonymous):

thank you.

OpenStudy (zpupster):

do you want to see addition?????????